OpenStudy (anonymous):

The probabilities of the top-selling menu item of the family restaurants in four towns being vegetarian and different types of non-vegetarian sandwiches are given in the table. Town Vegetable Sandwich Non-vegetarian Chicken Sandwich Non-vegetarian Ham Sandwich Non-vegetarian Bacon Sandwich Non-vegetarian Tuna Sandwich Town 1 6.63% 5.42% 7.05% 6.19% 5.91% Town 2 6.84% 7.11% 6.89% 5.69% 6.58% Town 3 5.97% 6.63% 5.72% 5.86% 5.79% Town 4 6.54% 5.86% 6.24% 6.24% 6.33% Total 6.24% 5.77% 6.03% 5.92% 6.27%% If a family restaurant in Town 2 has a non-vege

OpenStudy (anonymous):

Hi BSP!!!!!!!!!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

any answers?

OpenStudy (anonymous):

If a family restaurant in Town 2 has a non-vegetarian sandwich as its top-selling menu item, what is the probability that it is a tuna sandwich? 21.7% 25.0% 27.4% 62.7% 65.8%

OpenStudy (anonymous):

ummmmmmm @paki @SolomonZelman

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

Town/ Vegetable Sandwich/ Non-vegetarian Chicken Sandwich/ Non-vegetarian Ham Sandwich/ Non-vegetarian Bacon Sandwich/ Non-vegetarian Tuna Sandwich thats the order

OpenStudy (anonymous):

this problem sucks

OpenStudy (anonymous):

what grade are you in?

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

y

OpenStudy (anonymous):

im in 7th

OpenStudy (anonymous):

and im in honors so some of this stuff is easy to me but this one

OpenStudy (anonymous):

this one is evil

OpenStudy (mathmate):

@bsp50150 It's about conditional probability, which you probably just learned.

OpenStudy (mathmate):

Hint: Focus on town 2: <------------ non-veg-----------> vege tuna Town 2 6.84% 7.11% 6.89% 5.69% 6.58% Total of all sandwiches = 33.11% (\(\Omega\)) total of all non-vege = 26.27 (N) % of tuna = 6.58 (T) Probability it's tuna(T) sandwich given that it's a non-vege(N) is given by the conditional probability P(T|N)=P(T\(\cap\)N)/P(N)